OpenStudy (anonymous):

Jose left the White House and drove toward the recycling plant at an average speed of 40 km/h. Rob left some time later driving in the same direction at an average speed of 48 km/h. After driving for five hours Rob caught up with Jose. How long did Jose drive before Rob caught up?

OpenStudy (watchmath):

Suppose Jose left y hour earlier than Rob After 5 hours here are the distance traveled by the two Jose's distance= 40y+40(5)=200+40y Rob's distance = 48(5)=240 The two distances are equal 200+40y=240 40y=40 y=1 So Jose drove 1 hour earlier. It means when Rob caught up , Jose already drove for 5+1=6 hours.